OpenStudy (anonymous):

Can some one please explain this example question to me Part 1: If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (hba):

For part 1 If such situations occur there is no solution For part 2 There are 2 methods 1) Substitution 2)Elimination

OpenStudy (hba):

@mikala1

OpenStudy (anonymous):

yess

OpenStudy (hba):

Which one do you want to do Sub or Elim ?

OpenStudy (anonymous):

.. what is that

OpenStudy (hba):

A method to solve the equations.

OpenStudy (anonymous):

oo that thing i think it is sub

OpenStudy (anonymous):

am i correct

OpenStudy (hba):

No you can use 1 method Either Substitution or Elimination ? It is up to you.

OpenStudy (hba):

Which one do you want me to explain ?

OpenStudy (anonymous):

sub plz

OpenStudy (hba):

But Elimination is very easy

OpenStudy (anonymous):

ok then elim

OpenStudy (hba):

3y − 2x = 11 y + 2x = 9 ----------- ----------- Add them up ?

OpenStudy (hba):

@mikala1 ADD THEM UP !!

OpenStudy (anonymous):

where do i start .... i have nbo clue how to do this

OpenStudy (anonymous):

is it 3y -4x=20

OpenStudy (anonymous):

2x = 3y - 11 so y + (3y - 11) = 9 4y = 20 y = 5 5 + 2x = 9 2x = 4 x=2 So x=2 and y=5.

OpenStudy (hba):

You can check if your answer is correct by putting x=2 and y=5 in your equations and checking if, L.H.S=R.H.S if LHS=RHS then your answer is correct,If not then i am sorry your answer is incorrect

OpenStudy (anonymous):

lol its correct thanks youuu

OpenStudy (hba):

You are welcome :)

OpenStudy (unklerhaukus):

part 1 4 = 4 has infinite solutions

OpenStudy (hba):

Really wouldn't it be no solution ?

OpenStudy (unklerhaukus):

any value of x works

OpenStudy (hba):

hmm :/

OpenStudy (unklerhaukus):

4=3 has no solutions

OpenStudy (anonymous):

ik i got that one lol its infinaty

OpenStudy (hba):

@UnkleRhaukus Thanks for telling. My bad :/